# Flux-balance laws for spinning bodies under the gravitational self-force

### v1

@@ IV.1 Local flux-balance laws
We then replace the factors of $h^{R}_{a^{\prime}b^{\prime}}$ with integrals over delta functions, so that using Eq. (77),
@@
$\left\langle\delta\dot{\Gamma}^{A}(\tau^{\prime})\right\rangle_{\tau^{\prime}}=\lim_{\Delta\tau\to\infty}\frac{\Pi^{AB}}{C_{g}\Delta\tau}\int_{\mathcal{V}(\tau,\Delta\tau;r)}h^{R}_{a^{\prime}b^{\prime}}\nabla_{B}\boldsymbol{T}^{a^{\prime}b^{\prime}},$ (98)
@@
$\left\langle\delta\dot{\Gamma}^{A}(\tau^{\prime})\right\rangle_{\tau^{\prime}}=\Pi^{AB}\mathcal{F}^{\mathcal{W}(r)}_{B}\{\boldsymbol{h}^{R},\boldsymbol{h}^{S}\}.$ (99)
@@
Now, by bilinearity, there will be a contribution to Eq. (90) coming from the same term that appears on the right-hand side of Eq. (99), except with $h^{R}_{ab}$ and $h^{S}_{ab}$ switched. We now compute this contribution. First, note that Eq. (60) implies that
@@
By using the Hamiltonian alignment condition and reversing the factorization above, we can write $\nabla_{A}h^{R}_{a^{\prime\prime}b^{\prime\prime}}(X)$ as $[\nabla_{\bar{A}}h^{R}_{a^{\prime\prime}b^{\prime\prime}}(\bar{X})]_{\bar{X}\to X}$, and then interchange the order of integrals as before, obtaining
$\begin{split}\mathcal{F}^{\mathcal{W}(r)}_{A}\{\boldsymbol{h}^{S},\boldsymbol{h}^{R}\}&=-\left\langle\left[\nabla_{\bar{A}}\int_{\mathcal{V}(\tau,\infty;r)}h^{R}_{a^{\prime\prime}b^{\prime\prime}}(\bar{X})\boldsymbol{\mathcal{T}}^{a^{\prime\prime}b^{\prime\prime}}(X,\tau^{\prime})\right]_{\bar{X}\to X}\right\rangle\\ &=-\left\langle\left[\nabla_{\bar{A}}\delta H\{\Upsilon_{\tau^{\prime}-\tau}(X),\bar{X}\}\right]_{\bar{X}\to X}\right\rangle,\end{split}$ (101)
where we have undone the remaining steps that appeared in deriving Eq. (99).
@@
$\begin{split}\left\langle\delta\dot{\Gamma}^{A}\right\rangle=\Pi^{AB}\bigg{\{}&\nabla_{B}\langle\delta H[\Upsilon_{\tau^{\prime}-\tau}(X),X]\rangle\\ &+\mathcal{F}^{\mathcal{W}(r)}_{B}\{\boldsymbol{h}^{S},\boldsymbol{h}^{R}\}\bigg{\}}.\end{split}$ (103)
@@
$\begin{split}\left\langle\delta\dot{\Gamma}^{A}\right\rangle=\frac{1}{2}\Pi^{AB}\bigg{\{}&\mathcal{F}^{\mathcal{W}(r)}_{B}\{\boldsymbol{h},\boldsymbol{h}\}-I^{SS}(r)\\ &+\nabla_{B}\langle H(X^{\prime},X)\rangle\bigg{\}},\end{split}$ (105)
@@
IV.2 Vanishing of the “divergent” contribution
@@
Evaluating $I^{SS}_{A}(r)$ involves the computation of an integral over a sphere of radius $r$ around $\gamma$. As such, determining the value of $I^{SS}_{A}(r)$ requires investigating the structure of the integrand as a function of the normal vector $n_{a}\equiv\nabla_{a}r$: if it contains an odd number of factors of $n_{a}$, then it must vanish.
To show that this is the case, we consider the parity structure of the various tensors which appear in this integrand: expanding such a tensor $\boldsymbol{Q}$ in powers of $r$, we have that
$\boldsymbol{Q}=\sum_{n=N}^{\infty}\frac{1}{r^{n}}\boldsymbol{Q}_{n}.$ (108)
@@
As examples, any analytic function near $\gamma$ must have even parity, from the Taylor expansion near $\gamma$; this is the case for the background metric, for example. However, $h^{S}_{ab}$, as can be seen from Eqs. (98-100) and (124) of [51], has odd parity in Lorenz gauge [^9] Note that this particular part of the calculation is gauge-dependent: in the “highly regular” gauges of [64], the parity structure is quite different, and in the usual “radiation” gauges used near null infinity and the horizon there are singularities which appear near the particle [66]. As we do not need to compute the metric perturbation except for asymptotically, we are free to merely specify abstract properties for the gauge in which we are working: near the particle, it should approach the Lorenz gauge, and near null infinity and the horizon, it should approach radiation gauge.. Similarly, the volume element $\boldsymbol{\epsilon}^{\mathcal{W}}$ on $\mathcal{W}(\tau,\Delta\tau;r)$ has an explicit factor of $n_{a}$ when compared to the spacetime volume element $\boldsymbol{\epsilon}$: [^10] The plus sign here is due to the fact that $r$ increases as one approaches the surface $\mathcal{W}(\tau,\Delta\tau;r)$ from the inside; see the discussion in Appendix B.2 of [35] to see why this is the orientation that is compatible with Stokes’ theorem.
$\boldsymbol{\epsilon}=\boldsymbol{n}\wedge\boldsymbol{\epsilon}^{\mathcal{W}}$ (109)
@@
From the examples given above, it therefore follows that the integrand of $I^{SS}_{A}(r)$ must be odd, as it is constructed from two copies of $h^{S}_{ab}$ and the factor of $n_{a}$ from pulling back the volume element to $\mathcal{W}(\tau,\Delta\tau;r)$, with everything else that appears being even. Since we are evaluating this integrand at an even power in its Taylor series expansion (zero), it follows that $I^{SS}_{A}(r)$ is of the form of an integral over the sphere of constant $r$ of an odd number of factors of $n_{a}$, which vanishes. This proves Eq. (107), which shows that Eq. (105) becomes Eq. (91).
@@ V Flux-balance laws in the Kerr spacetime
Next, note that the retarded metric perturbation which we are considering vanishes on the past horizon and on past null infinity. This is just a consequence of the fact that it is a field with retarded boundary conditions (that is, no incoming radiation at null infinity, and no outgoing radiation at the horizon). Our integral over $\mathcal{B}$ then becomes one over the future horizon and future null infinity, which recovers Eq. (1), where we define the following flux term $\mathcal{F}_{A}$:
$\begin{split}\mathcal{F}_{A}\equiv\lim_{\Delta w\to\infty}\frac{1}{16\pi C_{g}\Delta w}\bigg{[}&\int_{\mathscr{I}^{+}(\Delta w)}\boldsymbol{\omega}\{\boldsymbol{h},\nabla_{A}\boldsymbol{h}\}\\ &+\int_{\mathscr{H}^{+}(\Delta w)}\boldsymbol{\omega}\{\boldsymbol{h},\nabla_{A}\boldsymbol{h}\}\bigg{]},\end{split}$ (110)
where $\mathscr{I}^{+}(\Delta u)$ and $\mathscr{H}^{+}(\Delta v)$ are portions of future null infinity and the future horizon of length $\Delta u$ (in retarded time) and length $\Delta v$ (in ingoing Eddington-Finkelstein time), respectively. Note that this conserved current is defined in terms of metric perturbations, and so to obtain a more practical flux-balance law one should rewrite the symplectic current in terms of Teukolsky variables by metric reconstruction, as was done for example in [62].
@@ V.1 Non-spinning particles
$\{\vartheta^{\alpha},P_{\beta}\}=A^{\alpha}{}_{\beta},$ (118)
for some matrix $A^{\alpha}{}_{\beta}$ that is just given by $\partial P_{\beta}/\partial J_{\alpha}$. This set of coordinates is sufficient for the discussion in this paper.
@@
$\left\langle\delta\dot{P}_{\alpha}(\tau^{\prime})\right\rangle_{\tau^{\prime}}=-A^{\beta}{}_{\alpha}(\partial_{\vartheta^{\beta}})^{A}[\mathcal{F}_{A}+\nabla_{A}\langle\mathcal{H}(\tau^{\prime},X)\rangle_{\tau^{\prime}}].$ (122)
In order to derive Eq. (2), we then need only show that
$\frac{\partial}{\partial\vartheta^{\alpha}}\langle\mathcal{H}(X,\tau^{\prime})\rangle_{\tau^{\prime}}=0.$ (123)
This equation follows from an argument analogous to that in Secs. 2.3 and 3.2 of [1], at least in the case of non-resonant background orbits. There is another, more straightforward way of understanding this argument which we present below.
@@
we therefore find that only $\mathcal{H}_{\boldsymbol{n},0}$ can be non-zero. As such, we can write
$\mathcal{H}(X,\tau^{\prime})=\sum_{\boldsymbol{n}}\mathcal{H}_{\boldsymbol{n}}(P)e^{in_{\alpha}[\vartheta^{\alpha}+(\tau^{\prime}-\tau)\nu^{\alpha}(P)]};$ (127)
@@
$\langle\mathcal{H}(X,\tau^{\prime})\rangle_{\tau^{\prime}}=\mathcal{H}_{0}(P),$ (128)
from which Eq. (123) immediately follows.
@@
$\begin{split}\left\langle\frac{\mathrm{d}}{\mathrm{d}\tau^{\prime}}\right.&\left.\left[\delta\vartheta^{\alpha}(\tau^{\prime})-(\tau^{\prime}-\tau)\frac{\partial\nu^{\alpha}}{\partial P_{\beta}}\delta P_{\beta}(\tau^{\prime})\right]\right\rangle_{\tau^{\prime}}\\ &=A^{\alpha}{}_{\beta}\left[(\partial_{P_{\beta}})^{A}\mathcal{F}_{A}+\frac{\partial}{\partial P_{\beta}}\langle\mathcal{H}(\tau^{\prime},X)\rangle\right].\end{split}$ (129)
This equation, however, is not entirely correct. First, the left-hand side diverges: the quantity inside the average does not have a well-defined average, as one can show that it grows linearly with $\tau^{\prime}-\tau$. This is also reflected in the right-hand side as well: the flux term also diverges, as $\partial_{P^{\alpha}}h_{a^{\prime}b^{\prime}}(X)$ grows linearly with $\tau^{\prime}-\tau$, and so the assumption that $\nabla_{A}h_{a^{\prime}b^{\prime}}(X)$ stays finite no longer holds. While neither side of Eq. (1) is strictly well-defined when one takes the angle component, it may be possible to recover useful results by instead considering both sides in a power series in $\tau^{\prime}-\tau$, and equating terms of equal powers. As this begins to stretch the notion of what one can mean by a “flux-balance law”, we leave a more careful exploration of this possibility to future work.

### v4

@@ IV.1 Local flux-balance laws
We then replace the factors of $h^{R}_{a^{\prime}b^{\prime}}$ with integrals over delta functions, so that, by using Eq. (79),
@@
$\left\langle\delta\dot{\Gamma}^{A}(\tau^{\prime})\right\rangle_{\tau^{\prime}}=\lim_{\Delta\tau\to\infty}\frac{\Pi^{AB}}{C_{g}\Delta\tau}\int_{\mathcal{V}(\tau,\Delta\tau;r)}h^{R}_{a^{\prime}b^{\prime}}\nabla_{B}\boldsymbol{T}^{a^{\prime}b^{\prime}},\vspace{\belowdisplayskip}$ (99)
@@
$\left\langle\delta\dot{\Gamma}^{A}(\tau^{\prime})\right\rangle_{\tau^{\prime}}=\Pi^{AB}\mathcal{F}^{\mathcal{W}(r)}_{B}\{\boldsymbol{h}^{R},\boldsymbol{h}^{S}\}.\vspace{\belowdisplayskip}$ (100)
@@
Now, by bilinearity, there will be a contribution to the flux term in Eq. (92) coming from the same term that appears on the right-hand side of Eq. (100), except with $h^{R}_{ab}$ and $h^{S}_{ab}$ switched. We now compute this contribution. First, note that Eq. (61) implies that
@@
By using the Hamiltonian alignment condition and reversing the factorization above, we can write $\nabla_{A}h^{R}_{a^{\prime\prime}b^{\prime\prime}}(X)$ as $[\nabla_{\bar{A}}h^{R}_{a^{\prime\prime}b^{\prime\prime}}(\bar{X})]_{\bar{X}\to X}$, and then interchange the order of integrals as before, and we can undo the remaining steps that appeared in deriving Eq. (100), obtaining
$\begin{split}\mathcal{F}^{\mathcal{W}(r)}_{A}\{\boldsymbol{h}^{S},\boldsymbol{h}^{R}\}&=-\left\langle\left[\nabla_{\bar{A}}\int_{\mathcal{V}(\tau,\infty;r)}h^{R}_{a^{\prime\prime}b^{\prime\prime}}(\bar{X})\boldsymbol{\mathcal{T}}^{a^{\prime\prime}b^{\prime\prime}}(X,\tau^{\prime})\right]_{\bar{X}\to X}\right\rangle_{\tau^{\prime}}\\ &=-\left\langle\left[\nabla_{\bar{A}}\delta H\{\Upsilon_{\tau^{\prime}-\tau}(X),\bar{X}\}\right]_{\bar{X}\to X}\right\rangle_{\tau^{\prime}}.\end{split}$ (102)
@@
$\begin{split}\left\langle\delta\dot{\Gamma}^{A}(\tau^{\prime})\right\rangle_{\tau^{\prime}}=\Pi^{AB}\bigg{\{}&\nabla_{B}\langle\delta H[\Upsilon_{\tau^{\prime}-\tau}(X),X]\rangle\\ &+\mathcal{F}^{\mathcal{W}(r)}_{B}\{\boldsymbol{h}^{S},\boldsymbol{h}^{R}\}\bigg{\}}.\end{split}$ (104)
@@
$\begin{split}\left\langle\delta\dot{\Gamma}^{A}(\tau^{\prime})\right\rangle_{\tau^{\prime}}=\frac{1}{2}\Pi^{AB}\bigg{\{}&\mathcal{F}^{\mathcal{W}(r)}_{B}\{\boldsymbol{h},\boldsymbol{h}\}-I^{SS}(r)\\ &+\nabla_{B}\langle\delta H[\Upsilon_{\tau^{\prime}-\tau}(X),X]\rangle_{\tau^{\prime}}\bigg{\}},\end{split}$ (106)
@@
IV.2 Conservative Hamiltonian
We now relate the third term on the right-hand side of Eq. (106) to the conservative Hamiltonian which appears in [49, 50, 51]. First, we have that
$\begin{split}\delta H[\Upsilon_{\tau^{\prime}-\tau}(X),X]&=\delta H[\Upsilon_{\tau^{\prime}-\tau}(X),\Upsilon_{\tau^{\prime}-\tau}(X)]\\ &\equiv 2\mathscr{H}[\Upsilon_{\tau^{\prime}-\tau}(X)],\end{split}$ (108)
by the fact that $H(\varepsilon)$ does not depend on where on the trajectory its second argument lies. We call $\mathscr{H}$, which is a true Hamiltonian (in the sense that it only depends on one point in phase space) the “coincidence Hamiltonian”. An argument similar to that which is used to derive Eq. (99) then shows that the average of the coincidence Hamiltonian is given by
$\left\langle\mathscr{H}[\Upsilon_{\tau^{\prime}-\tau}(X)]\right\rangle_{\tau^{\prime}}=\lim_{\Delta\tau\to\infty}\frac{1}{2C_{g}\Delta\tau}\int_{\mathcal{V}(\tau,\Delta\tau;r)}h^{R}_{a^{\prime}b^{\prime}}\boldsymbol{T}^{a^{\prime}b^{\prime}},$ (109)
where for brevity we have dropped the dependence on $X$ in each term of the integral on the right-hand side.
We now consider how the regular field is sourced by the stress-energy tensor: it is given in terms of an integral against a “regular two-point function” $G^{R}_{a^{\prime}b^{\prime}a^{\prime\prime}b^{\prime\prime}}$ [64]:
$\begin{split}h^{R}_{a^{\prime}b^{\prime}}(X)&=\int_{\mathcal{V}(\tau,\infty;r)}G^{R}_{a^{\prime}b^{\prime}a^{\prime\prime}b^{\prime\prime}}\boldsymbol{T}^{a^{\prime\prime}b^{\prime\prime}}(X)\\ &=\int_{\mathcal{V}(\tau,\Delta\tau;r)}G^{R}_{a^{\prime}b^{\prime}a^{\prime\prime}b^{\prime\prime}}\boldsymbol{T}^{a^{\prime\prime}b^{\prime\prime}}(X)\\ &\hskip 11.00008pt+R^{R}_{a^{\prime}b^{\prime}}(X,\Delta\tau),\end{split}$ (110)
where $R^{R}_{a^{\prime}b^{\prime}}(X,\Delta\tau)$ is a “remainder” term that must go to zero in the limit $\Delta\tau\to\infty$. As such, we find that
$\left\langle\mathscr{H}[\Upsilon_{\tau^{\prime}-\tau}(X)]\right\rangle_{\tau^{\prime}}=\lim_{\Delta\tau\to\infty}\frac{1}{2C_{g}\Delta\tau}\int_{\mathcal{V}(\tau,\Delta\tau;r)}\int_{\mathcal{V}(\tau,\Delta\tau;r)}G^{R}_{a^{\prime}b^{\prime}a^{\prime\prime}b^{\prime\prime}}\boldsymbol{T}^{a^{\prime}b^{\prime}}(X)\boldsymbol{T}^{a^{\prime\prime}b^{\prime\prime}}(X).$ (111)
Since the integrand is symmetric in the two copies of the stress-energy tensor, it must therefore follow that we can replace $G^{R}_{a^{\prime}b^{\prime}a^{\prime\prime}b^{\prime\prime}}$ with
$G^{\rm symm.}_{a^{\prime}b^{\prime}a^{\prime\prime}b^{\prime\prime}}=\frac{1}{2}\left(G^{R}_{a^{\prime}b^{\prime}a^{\prime\prime}b^{\prime\prime}}+G^{R}_{a^{\prime\prime}b^{\prime\prime}a^{\prime}b^{\prime}}\right).$ (112)
From this two-point function, we can define a “symmetric” part of the regular metric perturbation $h^{R}_{ab}$ by
$\begin{split}h^{\rm symm.}_{a^{\prime}b^{\prime}}(X)&\equiv\int_{\mathcal{V}(\tau,\infty;r)}G^{\rm symm.}_{a^{\prime}b^{\prime}a^{\prime\prime}b^{\prime\prime}}\boldsymbol{T}^{a^{\prime\prime}b^{\prime\prime}}(X)\\ &=\int_{\mathcal{V}(\tau,\Delta\tau;r)}G^{\rm symm.}_{a^{\prime}b^{\prime}a^{\prime\prime}b^{\prime\prime}}\boldsymbol{T}^{a^{\prime\prime}b^{\prime\prime}}(X)\\ &\hskip 11.00008pt+R^{\rm symm.}_{a^{\prime}b^{\prime}}(X,\Delta\tau),\end{split}$ (113)
for a similar remainder $R^{\rm symm.}_{a^{\prime}b^{\prime}}(X,\Delta\tau)$. As such, we can replace $h^{R}_{ab}$ with $h^{\rm symm.}_{ab}$ in Eq. (108) directly, and so
$\left\langle\delta H[\Upsilon_{\tau^{\prime}-\tau}(X),X]\right\rangle_{\tau^{\prime}}=2\left\langle\mathcal{H}[\Upsilon_{\tau^{\prime}-\tau}(X)]\right\rangle_{\tau^{\prime}},$ (114)
where $\mathcal{H}(X)$ is given by $\mathscr{H}(X)$, but using $h^{\rm symm.}_{ab}$ instead of $h^{R}_{ab}$. As discussed in [49, 50, 51], this is the true Hamiltonian which appears in the conservative dynamics at first order, and so we call it the “conservative Hamiltonian”.
IV.3 Vanishing of the “divergent” contribution
@@
Evaluating $I^{SS}_{A}(r)$ involves the computation of an integral over a sphere of radius $r$ around $\gamma$. As such, determining the value of $I^{SS}_{A}(r)$ can be performed by investigating the structure of the integrand as a function of the normal vector $n_{a}\equiv\nabla_{a}r$: if it contains an odd number of factors of $n_{a}$, then it must vanish.
To show that this is the case, we consider the parity structure of the various tensors which appear in this integrand: expanding such a tensor $\boldsymbol{Q}$ in powers of $r$, we have that (for some integer $N$)
$\boldsymbol{Q}=\sum_{n=N}^{\infty}r^{n}\boldsymbol{Q}_{n}.$ (116)
@@
As examples, any analytic function near $\gamma$ must have even parity, from the Taylor expansion near $\gamma$; this is the case for the background metric, for example. However, $h^{S}_{ab}$, as can be seen from Eqs. (98-100) and (124) of [67], has odd parity in Lorenz gauge. [^13] Note that this particular part of the calculation is gauge-dependent: in the “highly regular” gauges of [80], the parity structure is quite different, and in the usual “radiation” gauges used near null infinity and the horizon there are singularities which appear near the particle [82]. As we do not need to compute the metric perturbation except for asymptotically, we are free to merely specify abstract properties for the gauge in which we are working: near the particle, it should approach the Lorenz gauge, and near null infinity and the horizon, it should approach radiation gauge. Similarly, the volume element $\boldsymbol{\epsilon}^{\mathcal{W}}$ on $\mathcal{W}(\tau,\Delta\tau;r)$ has an explicit factor of $n_{a}$ when compared to the spacetime volume element $\boldsymbol{\epsilon}$: [^14] The plus sign here is due to the fact that $r$ increases as one approaches the surface $\mathcal{W}(\tau,\Delta\tau;r)$ from the inside; see the discussion in Appendix B.2 of [45] to see why this is the orientation that is compatible with Stokes’ theorem.
$\boldsymbol{\epsilon}=\boldsymbol{n}\wedge\boldsymbol{\epsilon}^{\mathcal{W}},$ (117)
@@
From the examples given above, it therefore follows that the integrand of $I^{SS}_{A}(r)$ must be odd, as it is constructed from two copies of $h^{S}_{ab}$ and the factor of $n_{a}$ from pulling back the volume element to $\mathcal{W}(\tau,\Delta\tau;r)$, with everything else that appears being even. Since we are evaluating this integrand at an even power in its Taylor series expansion (zero), it follows that $I^{SS}_{A}(r)$ is of the form of an integral over the sphere of constant $r$ of an odd number of factors of $n_{a}$, which vanishes. This proves Eq. (115), which [along with Eq. (114)] shows that Eq. (106) becomes Eq. (93).
@@ V Flux-balance laws in the Kerr spacetime
For concreteness, we now define a coordinate $w$ such that
$w(\tau)=t(\tau)+h(r),$ (118)
where $t$ and $r$ are in the usual Boyer-Lindquist coordinates, and $t(\tau)$ indicates the value of $t$ such that the background worldline $\gamma$ has coordinate time $t$ at proper time $\tau$. We moreover suppose that $h(r)$ has the property that
$h(r)=\begin{cases}r^{*}&r\to r_{H}\\ -r^{*}&r\to\infty\end{cases},$ (119)
where $r_{H}$ is the location of the horizon and $r^{*}$ is the tortoise coordinate in Kerr, such that $w$ becomes $v$ (ingoing Eddington-Finkelstein time) at the horizon and $u$ (retarded time) at null infinity. With $w$ defined, we define our outer region $\mathcal{B}$ such that $\mathcal{B}_{\pm}$ are surfaces of constant $w(\tau_{\pm})$.
Next, note that the retarded metric perturbation which we are considering vanishes on the past horizon and on past null infinity. This is just a consequence of the fact that it is a field with retarded boundary conditions (that is, no incoming radiation at null infinity, and no outgoing radiation at the horizon). Our integral over $\mathcal{B}$ then becomes one over the future horizon and future null infinity. However, we need to write the flux integral in Eq. (91) in terms of $w$, which is a good coordinate at these surfaces: to do so, note that
$\Delta w\equiv w(\tau_{+})-w(\tau_{-})=\int_{\tau_{-}}^{\tau_{+}}z(\tau^{\prime})\mathrm{d}\tau^{\prime},$ (120)
where $z\equiv\mathrm{d}t/\mathrm{d}\tau$ is the redshift for the background worldline. As such, we recover Eq. (1), where we define the following flux term $\mathcal{F}_{A}$:
$\begin{split}\mathcal{F}_{A}\equiv\lim_{\Delta w\to\infty}\frac{\langle z(\tau^{\prime})\rangle_{\tau^{\prime}}}{16\pi C_{g}\Delta w}\bigg{[}&\int_{\mathscr{I}^{+}(\Delta w)}\boldsymbol{\omega}\{\boldsymbol{h},\nabla_{A}\boldsymbol{h}\}\\ &+\int_{\mathscr{H}^{+}(\Delta w)}\boldsymbol{\omega}\{\boldsymbol{h},\nabla_{A}\boldsymbol{h}\}\bigg{]},\end{split}$ (121)
where $\mathscr{I}^{+}(\Delta u)$ and $\mathscr{H}^{+}(\Delta v)$ are portions of future null infinity and the future horizon of length $\Delta u$ and length $\Delta v$, respectively. Note that this conserved current is defined in terms of metric perturbations, and so to obtain a more practical flux-balance law one should rewrite the symplectic current in terms of Teukolsky variables by metric reconstruction, as was done for example in [78]. A trick for circumventing such a calculation, using the results of [1], is laid out in [43].
@@ V.1 Non-spinning particles
$\{\vartheta^{\alpha},P_{\beta}\}=\frac{\partial P_{\beta}}{\partial J_{\alpha}}\equiv A^{\alpha}{}_{\beta},$ (129)
This set of coordinates is sufficient for the discussion in this paper.
@@
$\begin{split}\left\langle\delta\dot{P}_{\alpha}(\tau^{\prime})\right\rangle_{\tau^{\prime}}=-A^{\beta}{}_{\alpha}\bigg{\{}&(\partial_{\vartheta^{\beta}})^{A}\mathcal{F}_{A}\\ &+\frac{\partial\langle\mathcal{H}[\Upsilon_{\tau^{\prime}-\tau}(X)]\rangle_{\tau^{\prime}}}{\partial\vartheta^{\beta}}\bigg{\}}.\end{split}$ (133)
In order to derive Eq. (2), we then need only show that the second term on the right-hand side of this equation vanishes. This follows from an argument analogous to that in Secs. 2.3 and 3.2 of [1], at least in the case of non-resonant background orbits. There is another, more straightforward way of understanding this argument which we present below.
@@
we therefore find that only $H_{\boldsymbol{n},0}$ can be non-zero. As such, we can write
$\mathcal{H}[\Upsilon_{\tau^{\prime}-\tau}(X)]=\sum_{\boldsymbol{n}}\mathcal{H}_{\boldsymbol{n}}(P)e^{in_{\alpha}[\vartheta^{\alpha}+(\tau^{\prime}-\tau)\nu^{\alpha}(P)]};$ (137)
@@
$\langle\mathcal{H}[\Upsilon_{\tau^{\prime}-\tau}(X)]\rangle_{\tau^{\prime}}=\mathcal{H}_{0}(P),$ (138)
from which it follows that the second term in Eq. (133) vanishes, and so Eq. (2) holds.
@@
$\begin{split}\left\langle\frac{\mathrm{d}}{\mathrm{d}\tau^{\prime}}\right.&\left.\left[\delta\vartheta^{\alpha}(\tau^{\prime})-(\tau^{\prime}-\tau)\frac{\partial\nu^{\alpha}}{\partial P_{\beta}}\delta P_{\beta}(\tau^{\prime})\right]\right\rangle_{\tau^{\prime}}\\ &=A^{\alpha}{}_{\beta}\left\{(\partial_{P_{\beta}})^{A}\mathcal{F}_{A}+\frac{\partial}{\partial P_{\beta}}\langle\mathcal{H}[\Upsilon_{\tau^{\prime}-\tau}(X)]\rangle\right\}.\end{split}$ (139)
This equation, however, is not entirely correct. First, the left-hand side diverges: the quantity inside the average does not have a well-defined average, as one can show that it grows linearly with $\tau^{\prime}-\tau$. This is also reflected in the right-hand side as well: the flux term also diverges, as $\partial_{P^{\alpha}}h_{a^{\prime}b^{\prime}}(X)$ grows linearly with $\tau^{\prime}-\tau$, and so the assumption that $\nabla_{A}h_{a^{\prime}b^{\prime}}(X)$ stays finite no longer holds.
There are two possible ways around this issue. First, note that, in general, there can exist points in phase space where $A^{\alpha}{}_{\gamma}\partial\nu^{\beta}/\partial P_{\gamma}$ is no longer invertible: this will occur where the mapping between action variables and frequencies is no longer injective; that is, along isofrequency curves. As such, there will exist at least one dual vector $V_{\alpha}$ such that
$V_{\alpha}A^{\alpha}{}_{\beta}\frac{\partial\nu^{\gamma}}{\partial P_{\beta}}=0,$ (140)
and so contracting $V_{\alpha}$ into both sides of Eq. (139) would give an expression which is well-defined, although it would not be a true flux-balance law, as it would still involve the local self-force through the conservative Hamiltonian. As another option, it may be possible to recover useful results by generally considering both sides of Eq. (139) in a power series in $\tau^{\prime}-\tau$, and equating terms of equal powers. As this begins to stretch the notion of what one can mean by a “flux-balance law”, we leave a more careful exploration of this possibility to future work.